9x+8y=-37
x=-2-3y
solve by substitution method

Answer by shellyhotgirl(2)

9x+8y=-37
x=-2-3y
solve by substitution method
use the first equation 9x+8y=-37 and substiute x in this equation with the second one
9(-2-3y)+8y=-37
-18-27y+8y=-37 multiply by 9
-27y+8y=-19 add 18 to each side
-19y=-19 add like terms and then divde to get y by itself
y=1
Now that you have y=1 plug it into the second equation to find what x is look something like
x=-2-3(1)